Calculat the ionization energy of hydrogen atom using bohrs theory
1
Expert's answer
2020-07-20T06:21:49-0400

ionization energy is the energy Required to ionise electron( Remove electron ) from outer most orbit of isolated gaseous atom .


For Hydrogen Atom Outer Most Orbir is 1st orbit or ground state .


So , Enegry Required to Remove Electron From nth orbit of hydrogen atom = 13.6 / n2 ev....


So , Enegry Required to Remove Electron From 1st orbit of hydrogen atom = 13.6 ev.




ionization energy of hydrogen atom = 13.6 ev .( Answer ) .




ev is the electron volt


1 ev is 1.6 *10-19 j .


So, ionization energy of hydrogen atom = 13.6 * 1.6 * 10-19j = 2.176 * 10-18 j. ( Answer ) .